(1 point) Solve the given initial value problem
y′=5+e^(y−5x+4)

y(0)=−4

The solution in the implicit form is F(x,y)=1,, where F(x,y)=
Answer #1

Answer: Given initial value problem is y= 5+ e 9-5074 Lie dy y 1 = y5344. e 5+ -> 0 ch put y-58+4=u hoy mes +5 substituting